OpenStudy (anonymous):

A college football coach wants to know if there is a correlation between his players leg strength and the time it takes for them to sprint 40 yards

OpenStudy (anonymous):

A college football coach wants to know if there is a correlation between his players leg strength and the time it takes for them to sprint 40 yard. He sets up the following test records the data. Every day for a week he counts how many times each player can leg press 350 pounds. The following week he has each player sprint 40 yards every day. Table shows the average number of leg press repetitions and the average 40 yard dash time (in seconds) for seven randomly selected players. What is the equation of the line of best fit? How many seconds should he expect a player to take to run 40 yards if that player can do 22 leg-press repetitions? Round any values to the nearest 10th if necessary

OpenStudy (anonymous):

Leg press (reps) 15, 18, 8, 30, 26, 12,21 40- yard dash: 5.2, 6.3, 6.8, 8.2, 8.0, 5.3, 5.9

OpenStudy (anonymous):

@directrix

Directrix (directrix):

Use the online calculator for this. Go over to: http://www.mathportal.org/calculators/statistics-calculator/correlation-and-regression-calculator.php Enter the data as instructed. Then, post here what you got for the line. Then, I'll check it. @Ella_Bear_

OpenStudy (anonymous):

@directrix y=4.536+0.107•x

Directrix (directrix):

Okay, let me see what I get.

OpenStudy (anonymous):

Okay

Directrix (directrix):

Here's what I got. We have one more thing to do after this.

Directrix (directrix):

We got the same regression line.

Directrix (directrix):

How many seconds should he expect a player to take to run 40 yards if that player can do 22 leg-press repetitions? Round any values to the nearest 10th if necessary

Directrix (directrix):

y=4.536+0.107•x y=4.536+0.107•22 Okay, crank this out for y and round to the nearest tenth.

OpenStudy (anonymous):

100?

Directrix (directrix):

I wonder if this is also supposed to be rounded to the nearest tenth. y=4.536+0.107•x y = 4.5 + .1*x Let me check your work.

Directrix (directrix):

I don't see how you got 100 here: y=4.536+0.107•22 y = 6.89 which is 6.9 to the nearest tenth.

OpenStudy (anonymous):

I added the 2 numbers together and multiplied by 22.

Directrix (directrix):

You did this: y=(4.536+0.107)•22 That is not the question. Only the .107 is multiplied by 22

Directrix (directrix):

y = 6.89 which is 6.9 to the nearest tenth.